A body travels for 15 sec starting from rest with constant acceleration. If it travels distances S1, S2, S3 in first five seconds, second five seconds, & next five seconds resp. the relation B/w S1,S2,S3 is ???

Grade:12th pass

1 Answers

Shobhit Varshney IIT Roorkee
9 years ago
Dear student,

Thanks for your post.
Let ‘v’ denotes speed and ‘a’ denotes acceleration.

At t=0, v0= 0
At t=5, v5= 5a
At t=10, v10= 10a

For interval-1, (t=0 to t=5)
S1=v0*5 + (1/2)*a*52= 12.5a

For interval-2, (t=5 to t=10)
S2= v5*5 + (1/2)*a*52= 37.5a

For interval-3, (t=10 to t=15)
S3= v10*5 + (1/2)*a*52= 62.5a

Therefore, S1: S2: S3= 1:3:5

thanks.